OpenStudy (anonymous):

What is the equation of the parabola in vertex form? x^2+6x-y+3=0 a. (x+9)=(y-3)^2 b. (x+6)=(y+6)^2 c. (x+3)^2=(y+6) d. (x+36)=(y+33)^2

OpenStudy (anonymous):

y=x^2+6x+3 y=x^2+6x+9-9+3 y=(x+3)^2-6

OpenStudy (anonymous):

c. (x+3)^2=(y+6)

OpenStudy (anonymous):

i suppose c